I have a word problem. I need help setting up the equation for problem.
110 rail cars make up a train. 75 rail cars are loaded with an average of 11.6% protein per car. What percent protein needs to be loaded on the remaining 35 cars to make the whole 110 car train 12% protein?
Thanks: Jon from Oklahoma

Answer by lwsshak3(6513)

I have a word problem. I need help setting up the equation for problem.
110 rail cars make up a train. 75 rail cars are loaded with an average of 11.6% protein per car. What percent protein needs to be loaded on the remaining 35 cars to make the whole 110 car train 12% protein?
**
let x=% protein needs to be loaded on 35 cars
11.6%*75+x%*35=12%*110
8.7+35x=13.2
35x=13.2-8.7=4.5
x=4.5/35
x≈.129 or 12.9%
% protein needs to be loaded on 35 cars=12.9%